# RRWNet: Recursive Refinement Network for Effective Retinal Artery/Vein Segmentation and Classification

### v4

@@ Abstract
The caliber and configuration of retinal blood vessels serve as important biomarkers for various diseases and medical conditions. A thorough analysis of the retinal vasculature requires the segmentation of the blood vessels and their classification into arteries and veins, typically performed on color fundus images obtained by retinography. However, manually performing these tasks is labor-intensive and prone to human error. While several automated methods have been proposed to address this task, the current state of art faces challenges due to manifest classification errors affecting the topological consistency of segmentation maps. In this work, we introduce RRWNet, a novel end-to-end deep learning framework that addresses this limitation. The framework consists of a fully convolutional neural network that recursively refines semantic segmentation maps, correcting manifest classification errors and thus improving topological consistency. In particular, RRWNet is composed of two specialized subnetworks: a Base subnetwork that generates base segmentation maps from the input images, and a Recursive Refinement subnetwork that iteratively and recursively improves these maps. Evaluation on three different public datasets demonstrates the state-of-the-art performance of the proposed method, yielding more topologically consistent segmentation maps with fewer manifest classification errors than existing approaches. In addition, the Recursive Refinement module within RRWNet proves effective in post-processing segmentation maps from other methods, further demonstrating its potential. The model code, weights, and predictions are publicly available at https://github.com/j-morano/rrwnet.
@@ 1 Introduction
Figure 1: The proposed framework, RRWNet, here applied for the segmentation and classification of retinal arteries and veins, consists of a W-shaped fully convolutional neural network consisting of two subnetworks. The output of the first subnetwork (Base) is iteratively refined by the second (Recursive Refinement) through a recursive mechanism.
@@
Irrespective of the chosen approach, state-of-the-art FCNN-based methods consistently encounter the challenge of manifest classification errors. These errors appear as regions where the predicted class contradicts the expected topological configuration of the target structures being segmented and classified. In the context of A/V segmentation and classification, these errors translate to unreasonably misclassified segments within predominantly correctly segmented vessels, as exemplified in Fig. 2.
Figure 2: Examples of manifest classification errors produced by a state-of-the-art FCNN-based method Morano et al., [2021]. (1-3) While most of the vessel is classified as artery, the model misclassifies the last part as vein. (4-6) The model often confuses the classification of vessels in crossing areas. These errors are easily detected by a human observer because they are inconsistent with the overall structure of the vascular tree, hence the term “manifest”.
@@
These errors arise from the propensity of FCNN-based models to classify vessels based on local characteristics of the input image, neglecting the global structural context of the vascular tree. To mitigate these errors, several approaches have been proposed Girard et al., [2019], Kang et al., [2020], Chen et al., [2022], Hu et al., [2024]. Some methods Girard et al., [2019], Kang et al., [2020] employ ad hoc post-processing techniques based on graph propagation. Alternatively, other methods Chen et al., [2022], Hu et al., [2024] have proposed combining standard pixel-wise segmentation losses with adversarial losses Goodfellow et al., [2014] and custom-designed losses focused on specific characteristics of the predicted maps, such as topological consistency. This combined approach aims to guide the model towards generating more topologically consistent segmentations. Despite their contributions, these approaches exhibit limitations in their applicability beyond the specific task of blood vessel segmentation and classification (i.e., limited generalizability), and their overall effectiveness in mitigating manifest classification errors remains limited.
@@
In this work, we introduce RRWNet (Fig. 1), a novel end-to-end deep learning framework for semantic segmentation that addresses the challenge of manifest classification errors in A/V segmentation and classification. The proposed framework defines a recursive FCNN consisting of two specialized subnetworks: a Base subnetwork, which receives the input image and produces base segmentation maps, and a Recursive Refinement (RR) subnetwork, which receives the base segmentation maps and iteratively refines them, correcting manifest classification errors. Extensive evaluation on multiple publicly available A/V segmentation and classification datasets demonstrates that the proposed RRWNet achieves state-of-the-art performance. Specifically, it produces segmentation maps that exhibit superior classification accuracy and topological consistency, with a lower prevalence of manifest classification errors compared to existing methods. Additionally, our experiments showcase the versatility of RRWNet by demonstrating the effectiveness of the RR subnetwork as a standalone post-processing technique. This method significantly enhances the classification accuracy and topological consistency of segmentation maps produced by other methods. For the sake of reproducibility, the source code, pre-trained model weights, and predicted results associated with RRWNet are publicly available on GitHub: https://github.com/j-morano/rrwnet.
@@ 2.1 Vessel segmentation and classification
The first methods for vessel segmentation on color fundus images were based on ad hoc image processing techniques Jiang and Mojon, [2003], Nain et al., [2004], Tolias and Panas, [1998], Staal et al., [2004] or traditional learning models such as artificial neural networks Sinthanayothin et al., [1999], Marín et al., [2011]. Today, FCNNs based on the U-Net architecture Ronneberger et al., [2015] have become the state of the art Jiang et al., [2018], Oliveira et al., [2018], Jin et al., [2019], Wang et al., [2021], Liu et al., 2023c , Tan et al., [2022], Liu et al., 2023a , Liu et al., 2023b .
@@
2.2 Reducing manifest A/V classification errors
Most of the aforementioned studies acknowledge the challenge of manifest classification errors, resulting from models favoring local input characteristics over the global structure of the vascular tree. Existing approaches to address this problem can be broadly categorized into two groups: ad hoc post-processing and learning based. In addition, most of the reviewed methods also incorporate generalistic or ad hoc preprocessing techniques that aim to improve the quality of the input images and to enhance the differences between arteries, veins, and the background.
Some common generalistic preprocessing techniques used in the state of the art include global contrast enhancement and local intensity normalization and CLAHE Pizer et al., [1987], Girard et al., [2019], Hemelings et al., [2019], Morano et al., [2021]. Differently, other works have proposed more specific preprocessing techniques focused on the task of A/V differentiation, such as Yi et al., [2023], who proposed a multi-channel module including Gabor and Canny filters.
@@ 2.2 Reducing manifest A/V classification errors
Conversely, learning-based methods Hu et al., [2024], Chen et al., [2022], Karlsson and Hardarson, [2022] aim to address manifest classification errors by incorporating additional losses and architectural modifications. Hu et al., [2024] introduced a multi-class point consistency module to generate artery and vein skeletons, which are then used to compute different consistency losses aimed at improving topological consistency and mitigating classification errors. For vessel segmentation, Tan et al., [2022] proposed the introduction of prior knowledge of the vascular topology and contrastive learning to improve the segmentation in lesion areas. Also focusing on topology, Chen et al., [2022] proposed a GAN-based method with a topological loss. In particular, they proposed to use a discriminator designed to rank, from lowest to highest, the topological connectivity of the ground truth, the predicted mask, and a randomly transformed mask. The ranking error is used as a loss to encourage the model to produce more topologically consistent segmentation maps. In addition, they proposed a new module that extracts the high-level topological features of the images to force the model to predict vascular segmentation maps with a topology similar to that of the manual annotations. Although these methods are interesting and fairly effective, they rely on task-specific mechanisms, and their performance remains limited. Finally, other works have also explored the use of active learning Fhima et al., [2024] and multimodal self-supervised pretraining Morano et al., [2020], Shi et al., [2024] to improve classification and segmentation performance in cross-dataset and few-shot settings. However, these methods require a much larger amount of labeled data than the aforementioned methods. For instance, Fhima et al., [2024] trained the model on a total of 240 images, while Morano et al., [2020] used 59 paired and semi-automatically registered retinography-fundus fluorescein angiography (FFA) pairs for pretraining, and Shi et al., [2024] acquired and pretrained the model on approximately 2000 triplets of retinography, venous phase FFA, and arterial phase FFA.
In contrast to these methods, we rely on a recursive refinement approach with two specialized subnetworks that implicitly leverage local and global information to iteratively correct manifest classification errors using limited annotated data.
@@ 2.3 Iterative refinement
A common approach consists of stacking multiple deep modules and training them in an end-to-end fashion Newell et al., [2016], Shen et al., [2017], Karlsson and Hardarson, [2022], Galdran et al., [2022], Tan et al., [2022]. This allows for modules specialized in solving the errors of the previous modules. For example, Newell et al., [2016] proposed a novel architecture composed of eight consecutive modules for pose estimation in natural images. During training, all modules receive supervision through comparison of their outputs with the ground truth. However, such methods often require a substantial number of parameters, leading to high memory and computational costs during both training and inference. To address this limitation, recent work has proposed the use of very lightweight encoder-decoder networks Karlsson and Hardarson, [2022], Galdran et al., [2022]. Using a custom architecture consisting of four U-Net-like networks, Karlsson and Hardarson, [2022] achieved state-of-the-art performance on various A/V segmentation and classification benchmarks. However, their approach requires extensive hyperparameter tuning to perform well, and the optimal hyperparameters were found to be dataset-specific. Their final model was achieved through an exhaustive search exploring various network configurations and loss functions. In particular, the authors experimented with different numbers of networks, layers, levels, and kernels in each network/layer, as well as different weights for the loss terms and the regularization parameters. Additionally, the generalization capabilities of the method have not been thoroughly evaluated. Galdran et al., [2022] proposed a similar approach for the same task using only two stacked U-Net-like networks, although with limited performance. In the same line, Tan et al., [2022] proposed an architecture with two cascaded lightweight U-Net-like networks for retinal vessel segmentation.
@@ 3 Contributions
    We propose RRWNet, a novel deep learning framework for recursively refining semantic segmentation maps to correct manifest classification errors. Our framework is the first to combine the advantages of module stacking and recursive refinement approaches by decomposing the network into two specialized parts: a Base subnetwork, which provides the initial segmentation maps, and a Recursive Refinement subnetwork, which iteratively refines them. Both subnetworks are trained jointly in an end-to-end manner.
@@
    We propose and publicly release an effective implementation of the proposed framework, based on FCNNs, for the automatic segmentation and classification of retinal vessels into arteries and veins in retinography images.
@@
    We demonstrate that RRWNet achieves state-of-the-art performance in A/V segmentation and classification on various public datasets (RITE, LES-AV, and HRF), showcasing the effectiveness of our framework. In particular, our method yields much more correct and topologically consistent segmentations than the state-of-the-art, opening up new applications in clinical practice.
@@
    Furthermore, we show that the Recursive Refinement subnetwork of RRWNet can be used as an effective and robust standalone post-processing technique, significantly improving the classification accuracy and the topological consistency of segmentation maps generated by other state-of-the-art methods. To the best of our knowledge, this is the first purely learning-based, stand-alone post-processing method for A/V classification.
@@ 4 Methods
Figure 3: Proposed approach for the segmentation and classification of arteries and veins. The input image is fed to the Base subnetwork, which produces coarse segmentation maps of arteries (A), veins (V) and blood vessels (BV). Next, the A/V segmentation maps are fed to the Recursive Refinement subnetwork, which recursively refines them for a certain number of iterations $K$.
@@
$\mathbf{\hat{y}}_{k}=f(\mathbf{x},\theta,k)=\begin{cases}f_{R}(f(\mathbf{x},\theta,k-1),\theta_{R})^{A,V}\oplus f_{B}(\mathbf{x},\theta_{B})^{BV}&\text{if}k>0\\ f_{B}(\mathbf{x},\theta_{B})&\text{if }k=0\end{cases}$ (1)
@@
${w_{k}=\begin{cases}1&\text{if }k=0\\ \frac{k}{Z}&\text{if }k>0\end{cases}}$ (4)
with the normalization factor $Z=\sum_{k=1}^{K}k=\frac{K(K+1)}{2}$.
@@
Figure 4: U-Net architecture of the two subnetworks. $N$ represents the number of base channels. $C$ represents the number of output channels. In our case, $N=64$.
@@ 5.1 Datasets
Figure 5: Examples of retinography images from different datasets and their corresponding A/V segmentation maps. (a-c) RITE. (d,e) LES-AV. (f-h) HRF. The segmentation maps are visualized as RGB images composed of the segmentation maps of arteries (Red channel), veins (Green channel) and vessels (Blue channel). This composition makes arteries appear magenta, veins appear cyan, crossings appear white (because they are both arteries and veins at the same time), and uncertain vessels appear blue (because they are not assigned to either artery or vein class, but only to vessel). * Annotations used for training and testing. † Second expert annotations.
@@ 5.2 Experiments
Hyperparameter search: Number of refinement steps
The number of refinement steps $K$ is an important hyperparameter of RRWNet. For this reason, we conducted a grid search on the RITE dataset to determine its optimal value. In particular, we evaluated the performance of RRWNet with the following values of $K$: 2, 3, 6, 8, and 11. Since we use 4-fold cross-validation for training the models, we join all the predictions on the validation set for the different 4 models and create a “joint validation set” with all these predictions. The evaluation was based on the average of the AUROC and AUPR for artery, vein, and BV segmentation, as well as the accuracy of A/V and BV/BG classification. The value that yielded the best overall performance according to these metrics was selected for subsequent experiments.
@@
We compared RRWNet with several state-of-the-art methods for A/V segmentation and classification on RITE, LES-AV and HRF datasets: Girard et al., [2019], Galdran et al., [2019], Ma et al., [2019], Hemelings et al., [2019], Kang et al., [2020], Morano et al., [2021], Chen et al., [2022], Galdran et al., [2022], Hatamizadeh et al., [2022], Karlsson and Hardarson, [2022], Yi et al., [2023], Hu et al., [2024]. Only methods trained on the datasets used in this work were included in our experiments. To comprehensively evaluate our approach, we used different evaluation protocols. In the first place, we employed the standard evaluation protocol in the field, focusing on the A/V classification accuracy for the intersection of predicted and GT vessels. In addition, we performed a more detailed comparison with recent state-of-the-art methods that either provided the predicted continuous segmentation maps or the necessary source code to obtain them for any of the datasets, since some of the evaluations cannot be performed without these maps. In particular, following Morano et al., [2021] and Chen et al., [2022], we compared the methods in terms of A/V classification performance for all vessel pixels in the GT, as well as as A/V segmentation performance using and several threshold-independent and topological metrics. The methods included in the comparison were Morano et al., [2021], Chen et al., [2022] [^6] https://github.com/o0t1ng0o/TW-GAN (accessed on 2023-12-16), Karlsson and Hardarson, [2022] [^7] https://github.com/robert-karlsson/av-segmentation (accessed on 2023-12-16), and Galdran et al., [2022]  [^8] https://github.com/agaldran/lwnet (accessed on 2023-12-16). In order to standardize the evaluation criteria, we did not perform any post-processing on the segmentation maps produced by these methods. Following previous work Galdran et al., [2019], Chen et al., [2022], Galdran et al., [2022], models were trained and tested separately on RITE and HRF, while LES-AV was used for cross-dataset evaluation (trained on RITE, tested on LES-AV). Statistical significance was assessed using the one-tailed Wilcoxon signed-rank test with sample-level bootstrapping.
@@
5.3 Evaluation metrics
A/V and BV/BG classification performances were evaluated using one-versus-one evaluation metrics (sensitivity, specificity, and accuracy), considering arteries and BV as the positive class, respectively. Only vessel pixels excluding crossings and uncertain vessels were involved in the calculation. While most prior works Girard et al., [2019], Hemelings et al., [2019], Galdran et al., [2022], Karlsson and Hardarson, [2022] only consider the intersection between predicted and GT vessels, this approach can yield misleading performance measures, especially for poor segmentations, and hinders standardized comparisons. Therefore, following recent works Morano et al., [2021], Chen et al., [2022], we also report classification performance for all vessel pixels in the GT, including those that were not detected as such by the models. In our state-of-the-art comparison, we report the classification performance for both scenarios, explicitly specifying the evaluation criteria used. Unless otherwise noted, results correspond to the test set of the corresponding dataset.
5.4 Training and evaluation details
We employed 4-fold cross-validation on the RITE and HRF training sets, dividing each fold into 80% for training and 20% for validation. The Adam optimizer Kingma and Ba, [2015] with a constant learning rate of $\alpha=1\times 10^{-4}$ and decay rates of $\beta_{1}=0.9$ and $\beta_{2}=0.999$ was used for training. Early stopping was applied after 200 epochs with no decrease in validation loss. The batch size is set to 1. In each fold, the best model was selected based on the validation loss. For the state-of-the-art comparison, the model with the lowest validation error was selected among the best models from the different folds.
We maintained the original splits of RITE and HRF for training and testing. RITE images were used at full resolution. HRF images were resized to 1024 pixels wide, but predicted segmentation maps were then upsampled to the original resolution for the evaluation, following Galdran et al., [2022]. Similarly, for LES-AV, we predicted at full resolution by feeding the model trained on RITE with LES-AV images resized to 576 pixels wide and then upsampling the predictions.
For INF and COR calculations, 1000 paths were used for RITE and 100 for HRF and LES-AV datasets, balancing computational cost with metric reliability. For bootstrapping, we used 5000 iterations and a sample size of $\frac{1}{2}$ of the total number of samples in the test set.
The methodology was implemented in Python 3 with PyTorch. The code, model weights, and test set predictions are available on GitHub: https://github.com/j-morano/rrwnet. The experiments were run on a server with dual AMD EPYC 7443 24-Core CPUs (1024GB of RAM) and one NVIDIA RTX A6000 GPU. Training RRWNet takes approximately 3 hours on this setup, while image segmentation takes under 0.1 seconds on the GPU and 6-8 seconds on the CPU.
6 Results and Discussion
6.1 Hyperparameter search: Number of refinement steps
Table 2: Impact of $K$ in RITE (joint validation set). Proposed RRWNet with different $K$: 2, 3, 6, 8, and 11. All “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. All values are percentages. Best results are highlighted in bold, and second best results are underlined.
@@ 6.1 Hyperparameter search: Number of refinement steps
6.2 Effect of the refinement module
Table 3 shows the evolution of the A/V and BV/BG classification performance of the selected RRWNet ($K=6$) over the refinement iterations on the RITE dataset.
Table 3: Evolution of the A/V and BV/BG classification performance of RRWNet (best model) in RITE over the refinement iterations. Since BV maps are not refined, BV/BG classification performance remains constant across iterations. In this case, all “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. All values are in percentages. The best values for each metric are highlighted in bold. The $p$-values of the Wilcoxon signed-rank test comparing the performance at $k=0$ and $k=6$ are also shown: **: $p<0.001$.
Iteration ($k$) A/V BV/BG
Acc. Sens. Spec. Acc. Sens. Spec. AUROC
0 89.62 86.39 92.26 96.29 80.16 98.61 98.50
1 93.42 92.22 94.41 = = = =
2 94.46 93.76 95.03 = = = =
3 94.78 94.19 95.26 = = = =
4 94.94 94.45 95.34 = = = =
5 94.92 94.42 95.34 = = = =
6 94.95** 94.46** 95.38** = = = =
The results show a progressive and, in the end, significant improvement in A/V classification performance over the refinement iterations. For instance, the accuracy increased significantly from 89.62% at $k=0$ to 94.95% (+5.33 pp) at $k=6$, with a corresponding increase in sensitivity and specificity. The improvement can also be observed at a glance in the segmentation maps. Fig. 6 displays the segmentation maps generated by the same model at each iteration $k$ for a sample image from the RITE dataset.
GT Base segmentation ($k=0$) $k=1$
Refer to caption Refer to caption Refer to caption
$k=2$ $k=3$ $k=4$
Refer to caption Refer to caption Refer to caption
$k=5$ $k=6$
Refer to caption Refer to caption
Figure 6: Effect of the refinement module of RRWNet on A/V classification. The iterative refinement approach progressively improves A/V classification. For this particular example, the initial accuracy of 86.86% ($k=0$) is improved to an accuracy of 88.89% ($k=6$). [RITE, image 05]
As evident in the figure, the RR module progressively improves the A/V classification over the iterations. Notably, in the base segmentation map ($k=0$), several misclassified vessels are visible. However, these errors are progressively reduced in subsequent iterations, resulting in a final segmentation map ($k=6$) with significantly fewer manifest classification errors and improved delineation of both arteries and veins.
6.3 Ablation study
Table 4: Ablation study in RITE. W-Net: 2 stacked U-Nets without recursive refinement ($K=1$), similar to Galdran et al., [2022]. RRU-Net: recursive refinement using a single U-Net module, similar to Mosinska et al., [2018]. RRWNetAll: proposed approach with recursive refinement ($K=6$) for arteries, veins, and BV. RRWNet: proposed approach with recursive refinement of arteries and veins only ($K=6$). The results are the average of the 4 models trained for each fold. All “vessel” pixels from the GT except crossings and unknown pixels are considered for the evaluation. A one-tailed Wilcoxon signed-rank test was performed to compare the results of the proposed RRWNet with the results of the best or second best model for each evaluation metric. *: $p<0.05$. Best results are highlighted in bold, and second best results are underlined.

### v5

@@ Abstract
The caliber and configuration of retinal blood vessels serve as important biomarkers for various diseases and medical conditions. A thorough analysis of the retinal vasculature requires the segmentation of the blood vessels and their classification into arteries and veins, typically performed on color fundus images obtained by retinography. However, manually performing these tasks is labor-intensive and prone to human error. While several automated methods have been proposed to address this task, the current state of art faces challenges due to manifest classification errors affecting the topological consistency of segmentation maps. In this work, we introduce RRWNet, a novel end-to-end deep learning framework that addresses this limitation. The framework consists of a fully convolutional neural network that recursively refines semantic segmentation maps, correcting manifest classification errors and thus improving topological consistency. In particular, RRWNet is composed of two specialized subnetworks: a Base subnetwork that generates base segmentation maps from the input images, and a Recursive Refinement subnetwork that iteratively and recursively improves these maps. Evaluation on three different public datasets demonstrates the state-of-the-art performance of the proposed method, yielding more topologically consistent segmentation maps with fewer manifest classification errors than existing approaches. In addition, the Recursive Refinement module within RRWNet proves effective in post-processing segmentation maps from other methods, further demonstrating its potential. The model code, weights, and predictions will be publicly available at https://github.com/j-morano/rrwnet.
@@ 1 Introduction
Fig. 1: The proposed framework, RRWNet, here applied for the segmentation and classification of retinal arteries and veins, consists of a W-shaped fully convolutional neural network consisting of two subnetworks. The output of the first subnetwork (Base) is iteratively refined by the second (Recursive Refinement) through a recursive mechanism.
@@
Irrespective of the chosen approach, state-of-the-art FCNN-based methods consistently encounter the challenge of manifest classification errors. These errors appear as regions where the predicted class contradicts the expected topological configuration of the target structures being segmented and classified. In the context of A/V segmentation and classification, these errors translate to unreasonably misclassified segments within predominantly correctly segmented vessels, as exemplified in Fig. 2. These errors arise from the propensity of FCNN-based models to classify vessels based on local characteristics of the input image, neglecting the global structural context of the vascular tree. To mitigate these errors, several approaches have been proposed Girard et al., [2019], Kang et al., [2020], Chen et al., [2022], Hu et al., [2024]. Some methods Girard et al., [2019], Kang et al., [2020] employ ad hoc post-processing techniques based on graph propagation. Alternatively, other methods Chen et al., [2022], Hu et al., [2024] have proposed combining standard pixel-wise segmentation losses with adversarial losses Goodfellow et al., [2014] and custom-designed losses focused on specific characteristics of the predicted maps, such as topological consistency. This combined approach aims to guide the model towards generating more topologically consistent segmentations. Despite their contributions, these approaches exhibit limitations in their applicability beyond the specific task of blood vessel segmentation and classification (i.e., limited generalizability), and their overall effectiveness in mitigating manifest classification errors remains limited.
Fig. 2: Examples of manifest classification errors produced by a state-of-the-art FCNN-based method Morano et al., [2021]. (1-3) While most of the vessel is classified as artery, the model misclassifies the last part as vein. (4-6) The model often confuses the classification of vessels in crossing areas. These errors are easily detected by a human observer because they are inconsistent with the overall structure of the vascular tree, hence the term “manifest”.
@@
In this work, we introduce RRWNet (Fig. 1), a novel end-to-end deep learning framework for semantic segmentation that address the challenge of manifest classification errors in A/V segmentation and classification. The proposed framework defines a recursive FCNN consisting of two specialized subnetworks: a Base subnetwork, which receives the input image and produces base segmentation maps, and a Recursive Refinement (RR) subnetwork, which receives the base segmentation maps and iteratively refines them, correcting manifest classification errors. Extensive evaluation on multiple publicly available A/V segmentation and classification datasets demonstrates that the proposed RRWNet achieves state-of-the-art performance. Specifically, it produces segmentation maps that exhibit superior classification accuracy and topological consistency, with a lower prevalence of manifest classification errors compared to existing methods. Additionally, our experiments showcase the versatility of RRWNet by demonstrating the effectiveness of the RR subnetwork as a standalone post-processing technique. This method significantly enhances the classification accuracy and topological consistency of segmentation maps produced by other methods. For the sake of reproducibility, the source code, pre-trained model weights, and predicted results associated with RRWNet will be made publicly available on GitHub: https://github.com/j-morano/rrwnet.
@@ 2.1 Vessel segmentation and classification
The first methods for vessel segmentation on color fundus images were based on ad hoc image processing techniques Jiang and Mojon, [2003], Nain et al., [2004], Tolias and Panas, [1998], Staal et al., [2004] or traditional learning models such as artificial neural networks Sinthanayothin et al., [1999], Marín et al., [2011]. Today, FCNNs based on the U-Net architecture Ronneberger et al., [2015] have become the state of the art Jiang et al., [2018], Oliveira et al., [2018], Jin et al., [2019], Wang et al., [2021], Liu et al., 2023c , Liu et al., 2023a , Liu et al., 2023b .
@@
2.2 Reducing manifest classification errors
Most of the aforementioned studies acknowledge the challenge of manifest classification errors, resulting from models favoring local input characteristics over the global structure of the vascular tree. Existing approaches to address this problem can be broadly categorized into two groups: ad hoc post-processing and learning based.
@@ 2.2 Reducing manifest classification errors
Conversely, learning-based methods Hu et al., [2024], Chen et al., [2022], Karlsson and Hardarson, [2022] aim to address manifest classification errors by incorporating additional losses and architectural modifications. Hu et al., [2024] introduced a multi-class point consistency module to generate artery and vein skeletons, which are then used to compute different consistency losses aimed at improving topological consistency and mitigating classification errors. Similarly, Chen et al., [2022] proposed a GAN-based method with a topological loss. In particular, they proposed to use a discriminator designed to rank, from lowest to highest, the topological connectivity of the ground truth, the predicted mask, and a randomly transformed mask. The ranking error is used as a loss to encourage the model to produce more topologically consistent segmentation maps. In addition, they proposed a new module that extracts the high-level topological features of the images to force the model to predict vascular segmentation maps with a topology similar to that of the manual annotations. Although these methods are interesting and fairly effective, they rely on task-specific mechanisms, and their performance remains limited.
In contrast to these methods, we rely on a recursive refinement approach with two specialized subnetworks that implicitly leverage local and global information to iteratively correct manifest classification errors.
@@ 2.3 Iterative refinement
A common approach consists of stacking multiple deep modules and training them in an end-to-end fashion Newell et al., [2016], Shen et al., [2017], Karlsson and Hardarson, [2022], Galdran et al., [2022]. This allows for modules specialized in solving the errors of the previous modules. For example, Newell et al., [2016] proposed a novel architecture composed of eight consecutive modules for pose estimation in natural images. During training, all modules receive supervision through comparison of their outputs with the ground truth. However, such methods often require a substantial number of parameters, leading to high memory and computational costs during both training and inference. To address this limitation, recent work has proposed the use of very lightweight encoder-decoder networks Karlsson and Hardarson, [2022], Galdran et al., [2022]. Using a custom architecture consisting of four U-Net-like networks, Karlsson and Hardarson, [2022] achieved state-of-the-art performance on various A/V segmentation and classification benchmarks. However, their approach requires extensive hyperparameter tuning to perform well, and the optimal hyperparameters were found to be dataset-specific. Their final model was achieved through an exhaustive search exploring various network configurations and loss functions. In particular, the authors experimented with different numbers of networks, layers, levels, and kernels in each network/layer, as well as different weights for the loss terms and the regularization parameters. Additionally, the generalization capabilities of the method have not been thoroughly evaluated. Galdran et al., [2022] proposed a similar approach for the same task using only two stacked U-Net-like networks, although with limited performance.
@@ 3 Contributions
    We propose RRWNet, a novel end-to-end deep learning framework for recursively refining semantic segmentation maps to correct manifest classification errors. Our framework is the first to combine the advantages of module stacking and recursive refinement approaches by decomposing the network into two specialized parts, a Base subnetwork and a Recursive Refinement subnetwork, which are trained jointly in an end-to-end manner.
@@
    We propose and publicly release a straightforward implementation of the proposed framework, based on FCNNs, for the automatic segmentation and classification of retinal vessels into arteries and veins in retinography images.
@@
    We demonstrate that RRWNet achieves state-of-the-art performance in A/V segmentation and classification on various public datasets (RITE, LES-AV, and HRF), showcasing the effectiveness of our framework.
@@
    Furthermore, we show that Recursive Refinement subnetwork of RRWNet can be used as an effective standalone post-processing technique, significantly improving the classification accuracy and the topological consistency of segmentation maps generated by other state-of-the-art methods.
@@ 4 Methods
Fig. 3: Proposed approach for the segmentation and classification of arteries and veins. The input image is fed to the Base subnetwork, which produces coarse segmentation maps of arteries (A), veins (V) and blood vessels (BV). Next, the A/V segmentation maps are fed to the Recursive Refinement subnetwork, which recursively refines them for a certain number of iterations $K$.
@@
$\mathbf{\hat{y}}_{k}$ $=f(\mathbf{x},\theta,k)$ (1)
=$=\begin{cases}f_{R}(f(\mathbf{x},\theta,k-1),\theta_{R})^{A,V}\oplus f_{B}(\mathbf{x},\theta_{B})^{BV},&k>0\\ f_{B}(\mathbf{x},\theta_{B}),&k=0\end{cases}$
@@
$w_{k}=\begin{cases}1,&k=0\\ \frac{1}{Z}\sum_{k=1}^{K}k\mathcal{L}_{\text{S}}\left(\mathbf{\hat{y}}_{k},\mathbf{y}_{k}\right),&k>0\end{cases}$ (4)
with the normalization factor $Z=\sum_{k=1}^{K}k$.
@@
Fig. 4: U-Net architecture of the two subnetworks. $N$ represents the number of base channels. $C$ represents the number of output channels. In our case, $N=64$.
@@ 5.1 Datasets
Fig. 5: Examples of retinography images from different datasets and their corresponding A/V segmentation maps. (a-d) RITE. (e,f) LES-AV. (f-h) HRF. The segmentation maps are visualized as RGB images composed of the segmentation maps of arteries (Red channel), veins (Green channel) and vessels (Blue channel). This composition makes arteries appear magenta, veins appear cyan, crossings appear white (because they are both arteries and veins at the same time), and uncertain vessels appear blue (because they are not assigned to either artery or vein class, but only to vessel). * Annotations used for training and testing. † Second expert annotations.
@@ 5.2 Experiments
Hyperparameter search
The number of refinement steps $K$ is an important hyperparameter of RRWNet. For this reason, we conducted a grid search on the RITE dataset (validation) to determine its optimal value. In particular, we evaluated the performance of RRWNet with the following values of $K$: 2, 3, 6, 8, and 11. The evaluation was based on the average of the AUROC and AUPR for artery, vein, and BV segmentation, as well as the accuracy of A/V and BV/BG classification. The value that yielded the best overall performance according to these metrics was selected for subsequent experiments.
@@
We compared RRWNet with several state-of-the-art methods for A/V segmentation and classification on RITE, LES-AV and HRF datasets: Girard et al., [2019], Galdran et al., [2019], Ma et al., [2019], Hemelings et al., [2019], Kang et al., [2020], Morano et al., [2021], Chen et al., [2022], Galdran et al., [2022], Hatamizadeh et al., [2022], Karlsson and Hardarson, [2022], Hu et al., [2024]. To comprehensively evaluate our approach, we used different evaluation protocols. In the first place, we employed the standard evaluation protocol in the field, focusing on the A/V classification accuracy for the intersection of predicted and GT vessels. Additionally, we performed a more in depth comparison with recent state-of-the-art methods that provided source code or continuous predicted segmentation maps for any of the datasets. In particular, following Morano et al., [2021] and Chen et al., [2022], we compared the methods in terms of A/V classification performance for all vessel pixels in the GT, as well as as A/V segmentation performance using and several threshold-independent and topological metrics. The methods included in the comparison were Morano et al., [2021], Chen et al., [2022] [^6] https://github.com/o0t1ng0o/TW-GAN (accessed on 2023-12-16), Karlsson and Hardarson, [2022] [^7] https://github.com/robert-karlsson/av-segmentation (accessed on 2023-12-16), and Galdran et al., [2022]  [^8] https://github.com/agaldran/lwnet (accessed on 2023-12-16). In order to standardize the evaluation criteria, we did not perform any post-processing on the segmentation maps produced by these methods. Following previous work Galdran et al., [2019], Chen et al., [2022], Galdran et al., [2022], models were trained and tested separately on RITE and HRF, while LES-AV was used for cross-dataset evaluation (trained on RITE, tested on LES-AV).
@@
Table 2: Impact of $K$ in RITE (validation). Proposed RRWNet with different $K$: 2, 3, 6, 8, and 11. Best results are highlighted in bold, and second best results are underlined. All values are percentages.
@@
5.3 Evaluation metrics
A/V and BV/BG classification performances were evaluated using one-versus-one evaluation metrics (sensitivity, specificity, and accuracy), considering arteries and BV as the positive class, respectively. Only vessel pixels excluding crossings and uncertain vessels were involved in the calculation. While most prior works Girard et al., [2019], Hemelings et al., [2019], Galdran et al., [2022], Karlsson and Hardarson, [2022] only consider the intersection between predicted and GT vessels, this approach can yield misleading performance measures, especially for poor segmentations, and hinders standardized comparisons. Therefore, following recent works Morano et al., [2021], Chen et al., [2022], we also report classification performance for all vessel pixels in the GT, including those that were not detected as such by the models. In our state-of-the-art comparison, we report the classification performance for both scenarios, explicitly specifying the evaluation criteria used.
Table 3: Ablation study in RITE. W-Net: 2 stacked U-Nets without recursive refinement ($K=1$), similar to Galdran et al., [2022]. RRU-Net: recursive refinement using a single U-Net module, similar to Mosinska et al., [2018]. RRWNetAll: proposed approach with recursive refinement ($K=6$) for arteries, veins, and BV. RRWNet: proposed approach with recursive refinement of arteries and veins only ($K=6$). A one-tailed Wilcoxon signed-rank test was performed to compare the results of the proposed RRWNet with the results of the best or second best model for each evaluation metric. *: $p<0.05$. Best results are highlighted in bold, and second best results are underlined.